The XYZ Corporation has $50 million in excess cash and no debt. The company expects to generate additional FCFs of $87 million per year in subsequent years. The XYZ Corporation has 25 million outstanding shares and has an unlevered cost of capital of 15 percent.

If the firm uses all excess cash to repurchase its shares in the open market, how many shares can the XYZ Corporation repurchase?

A. 985,360 shares

B. 1,245,625 shares

C. 1,541,369 shares

D. 1,984,127 shares
Answer #1

cum-dividend price per share of the XYZ Corporation
=(87/15%+50)/25
=25.20/share

Number of shares repurchased=50/25.20*10^6=1984127 shares